OpenStudy (anonymous):

Solve 5/12m + 2/3 = -3/4 for m

OpenStudy (jiteshmeghwal9):

\[\LARGE{\frac{15+24}{36m}=\frac{-3}{4}}\]

OpenStudy (jiteshmeghwal9):

now cross multiply it u will probably gt ur answer:)

OpenStudy (anonymous):

ok thanksss for the website and the help :D

OpenStudy (jiteshmeghwal9):

yw:)